if a=cosθ+isinθ ,b=cosΦ+isinΦ ,c=cosψ+isinΨ and a⁄b+b/c+c/a=1 then cos(θ-Φ)+cos(Φ-ψ)+cos(ψ-θ) equals _________ if e iθ =cosθ+isinθ then e 2micot-1p (pi+1/pi-1)equals


  1. if a=cosθ+isinθ ,b=cosΦ+isinΦ ,c=cosψ+isinΨ and a⁄b+b/c+c/a=1 then cos(θ-Φ)+cos(Φ-ψ)+cos(ψ-θ) equals _________

  2. if e =cosθ+isinθ then e2micot-1p(pi+1/pi-1)equals

Grade:12

1 Answers

Gokul Joshi AskiitiansExpert-IITK
13 years ago

Dear Khushboo,

You can write a=eθ, b=eΦ, c=eΨ

Now a/b+b/c+c/a = eθ-Φ+eΦ-Ψ+eΨ-θ = 1 (given). Taking the real part of both the sides you get cos(θ-Φ)+cos(Φ-ψ)+cos(ψ-θ)=1

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly. All the best.

Regards,
Askiitians Experts,
Gokul  Joshi